### v1

# Navigating Public Sentiment in the Circular Economy through Topic Modelling and Hyperparameter Optimisation

Abstract
To advance the circular economy (CE), it is crucial to gain insights into the evolution of public sentiments, cognitive pathways of the masses concerning circular products and digital technology, and recognise the primary concerns. To achieve this, we collected data related to the CE from diverse platforms including Twitter, Reddit, and The Guardian. This comprehensive data collection spanned across three distinct strata of the public: the general public, professionals, and official sources. Subsequently, we utilised three topic models on the collected data. Topic modelling represents a type of data-driven and machine learning approach for text mining, capable of automatically categorising a large number of documents into distinct semantic groups. Simultaneously, these groups are described by topics, and these topics can aid in understanding the semantic content of documents at a high level. However, the performance of topic modelling may vary depending on different hyperparameter values. Therefore, in this study, we proposed a framework for topic modelling with hyperparameter optimisation for CE and conducted a series of systematic experiments to ensure that topic models are set with appropriate hyperparameters and to gain insights into the correlations between the CE and public opinion based on well-established models. The results of this study indicate that concerns about sustainability and economic impact persist across all three datasets. Official sources demonstrate a higher level of engagement with the application and regulation of CE. To the best of our knowledge, this study is pioneering in investigating various levels of public opinions concerning CE through topic modelling with the exploration of hyperparameter optimisation.
keywords:
circular economy, topic modelling, machine learning, hyperparameter optimisation.
journal: Energy and AI
[…]
[label1]organization=School of Mathematical and Computer Sciences, Heriot-Watt University, city=Edinburgh, country=United Kingdom
[…]
[label2]organization=Faculty of Engineering, Imperial College London, city=London, country=United Kingdom
[…]
[label3]organization=Edinburgh Business School, Heriot-Watt University, city=Edinburgh, country=United Kingdom
[…]
[label4]organization=Faculty of Engineering and Physical Sciences, University of Surrey, city=Surrey, country=United Kingdom
[…]
[label5]organization=School of Management Science and Engineering, Nanjing University of Information Science and Technology, city=Nanjing, country=China
1 Introduction
The concept of the Circular Economy (CE) has received increased attention from both academics and practitioners in recent years. The primary objectives of the CE are commonly regarded as achieving the twin goals of economic prosperity and improved resources efficiency [1]. Despite this, the impacts of CE on social equity and public perceptions are often overlooked or inadequately addressed. The implementation of CE also plays a critical role in the transition towards net-zero emissions [2, 3, 4, 5, 6]. By reducing resource consumption and optimising waste management, CE contributes to lowering energy demand and decreasing greenhouse gas emissions [7]. This relationship is especially significant in fields such as energy and artificial intelligence, where the integration of advanced technologies and circular principles can lead to substantial improvements in resource efficiency and sustainability.
Moreover, public sentiment and understanding of CE remain notably limited. CE represents a paradigm shift from the traditional linear economy—characterised by the “take–make–waste” model [8] to one that promotes the principles of regeneration, material reuse, waste reduction, and pollution prevention. In a world grappling with pressing environmental challenges, public discourse on CE plays a pivotal role in shaping policies, business strategies, and societal perceptions of sustainable development. Therefore, it has become increasingly crucial to comprehend and interpret public sentiments concerning the CE. This urgency stems not only from the necessity to understand the intricacies of evolving public attitudes but also in the imperative to in-depth analyse the cognitive pathways that influence the public’s perception of recycled products within the broader framework of the CE [9].
The novel contributions of this paper are summarised as follows. Firstly, we conducted comprehensive and structured research based on our established framework. Our data collection efforts encompassed a wide range of sources, including posts from widely-used social media platforms such as Twitter and Reddit, where individuals express their views and thoughts on various topics, including the CE. Secondly, we analysed content from prominent news publishers like The Guardian, recognising their crucial role in disseminating information as well as shaping public discussions and influencing the topics of conversation. Thirdly, the study provides a comprehensive understanding of how the public perceives the CE and how these perceptions change over time. By drawing data from diverse sources and engaging with these varied segments of society, the study captures a comprehensive view on shift in public sentiment and perspectives related to the CE.
Throughout the process of our study, we meticulously designed and executed a series of systematic experiments with the primary goal of unravelling the intricate relationship between the CE and public sentiment. Our aim was to move beyond mere observation and dive into the underlying factors driving changes in sentiment. To achieve this, we employed topic modelling techniques [10]. To be more specific, we utilised three well-established and open-source topic modelling techniques for their reliability and effectiveness. These include models based on Latent Dirichlet Allocation (LDA) technology [11], Correlation Explanation (CorEx) [12], and BERTopic model [13], which leverages Large Language Model technique on topic modelling. Note that these advanced tools enabled us not only to identify topics related to the CE across various platforms but also to track the ever-evolving and dynamic aspects of public opinion towards CE.
Furthermore, this study explores the complexity of our three models. We have not only implemented training of different models across various datasets but have also fine-tuned the hyperparameters of each model (hyperparameter optimisation). The hyperparameter optimisation process employed a grid search strategy, and optimal hyperparameters were identified for different datasets based on Normalised Pointwise Mutual Information (NPMI) scores [14, 15]. This approach ensures the efficacy and robustness of our analysis. Through the hyperparameter optimisation of the three models, this research found that BERTopic performed best on both The Guardian and Reddit datasets. CorEx achieved the highest NPMI score on Twitter dataset. In Sections 2 and 3, we will discuss the related work and the proposed framework for this study, as well as the experimental settings, respectively. In Section 4, we will comprehensively discuss and analyse each optimal model’s generated topics related to the CE, as well as the evolution of public sentiment towards CE topics over the past few years.
In essence, our research work stands as a pioneering initiative in the exploration of public sentiment across all facets related to the CE. More importantly, we aim to provide valuable insights into the pathways and perspectives guiding us towards a more sustainable future for the CE. We recognise that public sentiment serves as a compass, guiding the trajectory of policies, corporate strategies, and societal attitudes. By shedding light on this intricate relationship, we aim to contribute to the global effort to build a more sustainable CE that benefits us all.
2 Related work
In this section, we thoroughly examine the extant research and approaches relevant to our study’s scope and objectives. Our study is primarily oriented towards discerning public concerns regarding pivotal aspects of the CE by employing advanced techniques such as topic modelling and the utilisation of large language models. To facilitate a structured and comprehensive review, we categorise the related work into six subsections, each dedicated to an essential facet of our research: Latent Dirichlet Allocation (LDA), Correlation Explanation (CorEx) Topic Model, BERTopic, Evaluation Metrics, and Circular Economy. This section is intended to provide readers with a profound understanding of the foundational research that underpins our investigation.
2.1 Latent Dirichlet Allocation (LDA)
LDA is a technique for discovering the underlying topic structure inherent in a corpus of text documents. LDA operates within the framework of unsupervised machine learning [16] and provides a way to convert inherently qualitative textual information into a quantitative format that can be used for rigorous statistical modelling by representing each document as a vector of word frequencies. LDA assumes that each document in a collection of documents can be conceptually understood as a combination of multiple underlying topics, each characterised by a distribution of multiple words [11].
Understanding the core concepts of LDA is essential for gaining insight into topic modelling. LDA is built upon two fundamental equations, denoted as Equation 1 and Equation 2. For the sake of clarity, we introduce a new variable, denoted as $Q$, which signifies the relationship between each word in each document and the associated topic. $P(W|\alpha,\beta)$ represents the likelihood probability [11] of the entire document collection. The parameter $\alpha$ plays a pivotal role in controlling the mixture of topics within each document, while $\beta$ governs the distribution of words within each topic, determining which words are associated with specific topics. $z_{m,n}$ denotes the topic of the n-th word in the m-th document, $w_{m,n}$ signifies the n-th word in the m-th document, and $\beta_{z_{m,n}}$ represents the vocabulary distribution of topic $z_{m,n}$. In this study, we will not explain the mathematical derivation of LDA in depth. With core Equation 1 and Equation 2 of LDA, we aim is to provide readers with an intuitive understanding of LDA’s operational principles and its robust foundation as a powerful probabilistic model [11].
[…]
$Q=\prod_{n=1}^{N}\sum_{z_{m,n}}P(z_{m,n}|\theta_{m})P(w_{m,n}|\beta_{z_{m,n}}).$ (2)
Mahanty et al. [17] conducted an in-depth investigation into the evolution of the CE concept within the realm of academic articles. Their approach involved the utilisation of a straightforward and intuitive LDA model to compute probability distributions after topic generation. This LDA model was leveraged to analyse and compare the topics within each year. These probability distributions were then meticulously examined to discern prevalent trends in topics across the years. It is crucial to underscore that their analysis of LDA revealed an inherent limitation in the model’s capacity to accurately capture the temporal patterns embedded within a corpus of academic texts. Recognising this constraint, they thoroughly explored alternative models tailored to elucidate the dynamics of temporal changes within their academic text dataset.
Among the models designed to address the temporal dimension, the Topics Over Time (TOT) model [18] stands out for its distinctive feature of employing beta-distributed topics, deviating from the convention of Gaussian-distributed topics typically found in Dynamic Topic Models (DTM) [19]. It is noteworthy that the applicability of the TOT model is limited by its constraint of permitting only a point, allowing for one instance of rising and falling trends. This constraint imposes a significant limitation on the model’s ability to capture intricate and multifaceted changes in topics, such as scenarios involving a sequence of topic frequency decline, upswing, and subsequent decline [20]. Our text data exhibits analogous characteristics. Given that our data primarily revolve around the collection of public sentiments and the meticulous analysis of textual content, we anticipate the emergence of fluctuations in topic distributions over time. Hence, we opt for an initial approach utilising the fundamental LDA model to conduct a probabilistic-based topic modelling of the CE. In the study, we employ the fundamental Bag-of-Words (BoW) model [21] as a foundational element in our LDA model. The BoW model focuses exclusively on the presence or absence of individual words within documents, without regard to the sequential order in which they appear. This approach treats phrases such as ‘The Britain loves the circular economy’ and ‘The circular economy loves the Britain’ as equivalent, as the model disregards the specific word order. By discovering the recurrent topics found in the corpus and assessing how common they are in each document, this allows us to establish a foundational understanding of CE topics within the corpus.
Subsequently, we extend our analysis by employing other models: Correlation Explanation (CorEx) Topic Model and BERTopic, offering distinct methodologies for topic modelling. This multi-pronged strategy facilitates a comprehensive exploration of the dynamics surrounding public sentiment and its evolving facets regarding the concept of CE, enabling a deeper comprehension of sentiment changes within the public sentiment.
2.2 Correlation Explanation (CorEx) Topic Model
Correlation Explanation (CorEx) is a semi-supervised [22] topic model aimed at uncovering latent structures and topics within data. In contrast to traditional LDA, CorEx leverages the principle of maximum correlation from information theory, seeking to identify a set of latent topics that maximise the mutual information between the data and the topics [12].
CorEx’s objective is to discover the hidden topics by maximising the mutual information $I(X;T)$ between the data variables $X$ and the latent topic variables $T$. Formally, it seeks to maximise:
$\max_{p(t|x)}I(X;T)=\max_{p(t|x)}\sum_{x,t}p(x,t)\log\frac{p(x,t)}{p(x)p(t)},$ (3)
where $p(x,t)$ is the joint distribution [23] of $X$ and $T$, and $p(x)$ and $p(t)$ are the marginal distributions [24] of $X$ and $T$, respectively. To compute $I(X;T)$, CorEx employs a KL-divergence-based approximation [25], defining a tractable lower bound:
$I(X;T)\geq\mathbb{E}_{p(t|x)}\left[\sum_{x,t}p(t|x)p(x)\log\frac{p(t|x)}{r(t)}\right],$ (4)
where $r(t)$ is an estimate of the marginal distribution of $T$. By maximising this lower bound, CorEx can learn $p(t|x)$, the distribution of topics $T$. CorEx’s strength lies in its semi-supervised nature, enabling the incorporation of prior knowledge to guide the topic discovery process, thereby yielding more interpretable and meaningful topics. In this research, we designated the words ‘reduce’, ‘reuse’, and ‘recycle’ from the 3Rs policy [26] as anchor words (prior knowledge) for the CorEx model, leveraging this prior knowledge to steer the model towards discovering topics pertinent to the circular economy. By employing these anchor words as known topic words, CorEx endeavours to uncover other highly correlated words and topics, better capturing the semantics of the circular economy.
2.3 BERTopic
BERTopic utilises the pre-trained language model BERT [27] to identify topics in text data. In the study, we employ BERTopic to gain insights into public sentiment regarding the circular economy. The BERTopic with the class-based c-TF-IDF [13] procedure can leverage contextual understanding of words and phrases, and identify topics that may not be apparent through traditional methods [13]. Alamsyah et al. [28] sheds light on the remarkable capabilities of BERTopic in the analysis of sentiment within textual data. In their comparative analysis, BERTopic stands out as an effective model, achieving a notable coherence score [28] of $0.67$ when applied to user comments. This performance is particularly impressive when considering that the text inputs are often concise, aligning with the short text data paradigm, as previously discussed in the Correlation Explanation (CorEx) Topic Model. However, they [13, 28] do not consider the performance of BERTopic on non-short texts (such as our The Guardian news data), nor does it examine whether BERTopic can maintain the same effect on different types of data. As we embark on the study, it is essential to acknowledge that the dataset will encompass text of varying lengths, including non-short text [29]. This diversity in textual data prompts a natural curiosity regarding the adaptability and performance of BERTopic in handling these more extended textual inputs.
Figure 1: The process of BERTopic topic modelling
The initial steps of BERTopic involve the basic operations of converting words from the English language to numeric or vector representations (Embedding Vector). As shown in Figure 1, this key process plays an important role in the BERTopic model to recognise human language. This transformation is used to encode the semantics of words into a format suitable for BERTopic model manipulation, thus facilitating our subsequent topic modelling. Uniform Manifold Approximation and Projection (UMAP) will transform the vector obtained in BERTopic into a lower dimensional representation, so as to facilitate BERTopic to understand the data more intuitively. UMAP first finds the nearest neighbours. We can specify the number of neighbours by adjusting the $n_{neighbors}$ hyper-parameter of UMAP. UMAP uses the form of constructing relations in higher dimensions. For $K$ adjacent points $n$ of a certain point $m$, there is a relationship of Equation 5. Where $d_{\min(m)}$ represents the Euclidean distance [30] from the point to its nearest neighbour. The UMAP algorithm calculates the distance between each point $m$ and its closest neighbour, represented as $d_{\min(m)}$. Subsequently, it employs a Binary Search algorithm as described by Williams [31] to determine the optimal value of $\sigma_{m}$ required for solving the conditional probability $p_{m|n}$. This iterative process continues until the joint probability [32] $p_{mn}$ which is defined in Equation 6, converges to the desired state.
$p_{m|n}=e^{-\sigma_{m}\cdot d(x_{m},x_{n})-d_{\min(m)}},$ (5)
$p_{mn}=p_{m|n}+p_{n|m}-p_{m|n}\cdot p_{n|m},$ (6)
$q_{mn}=\frac{1}{1+a(y_{m}-y_{n})^{2b}}.$ (7)
UMAP uses the representation of joint probability to represent high-latitude data. McInnes et al. [33] designed a specific family of curves $\frac{1}{1+a(y^{2b})}$ to model the probability of distance in low dimensions, and Equation 7 directly represents the relationship between points $m$ and $n$ in the low dimensions $q_{mn}$. Both $a$ and $b$ in Equation 7 are hyper-parameters in UMAP. In this step, UMAP solves for the hyper-parameters $a$ and $b$ using $min_{d}ist$ a representation of the minimum distance between points in low dimensions. Thus, we can obtain the initial low-dimensional matrix.
In the last step of the whole UMAP algorithm, UMAP will use the Loss function [34] to make the representation of high dimensional points close to the representation of low dimensional points to achieve the effect of dimensional reduction of high-dimensional data. And according to McInnes et al. [33], UMAP will use cross-entropy [35] as the Loss function instead of Kullback Leibler Divergence [36]. This design can reduce the dimensional of the data by preserving the global distance when the data is moved from the high dimensional to the low dimensional [33]. In this paper, we will keep the default setting of UMAP, and cross-Entropy will be used as the loss function of UMAP to realise the dimensionality reduction step of BERTopic. After reducing the data’s dimensional using UMAP, we can employ an appropriate approach (Hierarchical Density-Based Cluster Selection (HDBSCAN) [37] or K-Means [38, 39]) to cluster the data effectively. HDBSCAN is a density clustering algorithm capable of identifying clusters with different shapes and sizes without the need to pre-specifying the number of clusters. K-Means is a clustering algorithm that identifies $K$ clusters within a given dataset. It discovers $K$ distinct clusters, with the centre of each cluster calculated as the mean of the values contained within that cluster. The number of clusters, $K$, is specified by the user, and each cluster is described by its centroid, which is the central point of all the points in the cluster.
From the early experiments, we found that both HDBSCAN and K-Means aided BERTopic in better comprehending the clustering structure of the dimensionality-reduced data. These two methods proved valuable when dealing with more complex and noisy datasets. However, since ensuring the quality of each acquired data point in real-world scenarios is typically challenging and this paper will involve a comparative analysis of three models. Controlling the number of topics generated or clusters formed by different models can provide a clearer understanding of their performance on the same dataset. K-Means allows us to select the desired number of clusters and forces every point to belong to a cluster. Consequently, no outliers will be produced. Nevertheless, as mentioned in BERTopic [13], there are drawbacks. When we forcibly cluster every point, it implies that the cluster is likely to contain noise that may degrade the topic representation. Fortunately, we can leverage the built-in English stop word library in the model to remove unnecessary vocabulary, thereby enhancing the topic representation. All in all, K-Means clustering algorithm will be the choice for this research.
$\text{TF}_{i,j}=\frac{n_{i,j}}{\sum_{N}\cdot n_{i,j}},$ (8)
$\text{IDF}(i)=\log\left(\frac{N}{N_{i}+1}\right),$ (9)
$W(t,c)=\text{TF}_{t,c}\cdot\log(1+\frac{A}{t\cdot f_{t}}).$ (10)
BERTopic modifies the Term Frequency-Inverse document frequency (TF-IDF) algorithm [40]. As the last step of the model, it uses class-based TF-IDF (c-TF-IDF) [13] to extract the topic words. TF-IDF is a statistical technique, same as TF $\cdot$ IDF, wherein TF stands for term frequency, and IDF stands for inverse document frequency. In Equation 8, $i$ represents a specific term, $j$ corresponds to a particular document, and $N$ denotes the total number of words within the document. Equation 8 allows us to compute the term frequency (TF) of a term $i$, which can also be interpreted as the probability of the term $i$ occurring in the content $j$. The IDF for a term $i$ is computed by dividing $N$ by $N_{i}$ which represents the number of documents containing the term $i$ and taking the algorithm of the resulting quotient (Equation 9). However, in cases where term $i$ does not appear in any document, it could lead to a division by zero issue. To prevent this, it is a common practice to $+1$ to the numerator to ensure the formula remains solvable.
From Robertson’s research [40], we understand that TF-IDF is employed to assess the significance of words. When the term frequency (TF) of word $i$ is high in one text but low in other articles, it is deemed that this word $i$ possesses strong category-distinguishing characteristics and is suitable for classification. However, the basic structure of IDF (Inverse Document Frequency) does not take into account the semantic nuances of words or effectively capture word distribution, rendering the IDF’s methodology somewhat rudimentary [41]. Consequently, it struggles to address issues related to word polysemy and homonymy, and as a result, the accuracy of the TF-IDF algorithm is not particularly high, especially when working with pre-classified text datasets [41]. Fortunately, researchers have made advancements in text representations for classification. BERTopic leverages an adapted TF-IDF model, enabling the extraction of keywords based on classified text, which is facilitated through the c-TF-IDF approach [13, 41]. In Equation 10, within the context of the c-TF-IDF algorithm. $W(t,c)$ denotes the weighted score (importance degree) [13] of word $t$ in the category $c$. $f_{t}$ represents the frequency of word $t$ across all categories, and $A$ signifies the average word count in each category. A notable departure from the traditional TF-IDF method can be observed in the IDF formula (Equation 9).
Following the modification of the c-TF-IDF algorithm, it becomes apparent that as $f_{t}$ increases, the likelihood of the word $t$ being a common term rather than one specific to a particular category also rises. In our research, we aim to achieve results similar to those in Sun et al. [41]. BERTopic effectively addresses a scenario where the c-TF-IDF algorithm may inadvertently overlook keywords in similar texts, thereby enhancing the accuracy of topic modelling.
2.4 Evaluation Metrics
Several metrics have been suggested for assessing the effectiveness of topic modelling techniques. For example, researchers have assessed the performance of topic models by gauging their accuracy in information retrieval, as demonstrated in prior work [42]. Additionally, statistical approaches have been employed to evaluate a topic model’s predictive capability on unseen documents, typically quantified through perplexity calculations [10]. However, topic coherence metrics have emerged as a predominant and widely used metric for evaluating topic models in recent years [43]: For a given topic, we select the $N$ most relevant terms as representative words. We pair these $N$ representative words to form a set of bigrams $\frac{(N(N-1))}{2}$. For each bigram, denoted as $(\omega_{1},\omega_{2})$. For each bigram, we compute a score using the Normalised Pointwise Mutual Information (NPMI) [14, 15]. NPMI is a normalisation operation based on Pointwise Mutual Information (PMI) [15, 44]. PMI between two words $\omega_{1}$ and $\omega_{2}$ is defined as:
$\text{PMI}(\omega_{1},\omega_{2})=\log\frac{P(\omega_{1},\omega_{2})}{P(\omega_{1})P(\omega_{2})},$ (11)
where $P(\omega_{1},\omega_{2})$ is the probability of the two words co-occurring, and $P(\omega_{1})$ and $P(\omega_{2})$ are the individual probabilities of each word occurring. PMI measures the relationship between two words by comparing the observed co-occurrence probability with the expected probability if the words were independent. However, PMI is biased towards infrequent co-occurrences, which can lead to unintuitive results for topic coherence evaluation. To address this issue, the Normalised PMI (NPMI) is employed [15]:
$\text{NPMI}(\omega_{1},\omega_{2})=\frac{\text{PMI}(\omega_{1},\omega_{2})}{-\log P(\omega_{1},\omega_{2})},$ (12)
NPMI normalises PMI to range between -1 and 1, where 1 indicates perfect coherence, 0 indicates independence, and -1 indicates perfect incoherence.
We then take the average of the scores of all bigrams within a topic as the coherence score for that topic. Finally, we average the coherence scores across all topics to obtain the overall coherence score for each model.
2.5 Circular Economy
In recent years, the concept of the CE has become a major paradigm shift, emerging as a viable alternative to the conventional linear economic model. By focusing on the regenerative and ethical use of resources throughout their lifespan, CE aims to minimise waste and maximise resource efficiency. At its core are the principles of reducing, reusing, and recycling, synergistically creating an economic framework that is more ecologically friendly and sustainable [9].
One strand of research examined the public recognition and understanding of the concept of the CE from broader social perspective [45, 46]. Some authors focused on exploring consumers’ attitudes and purchase intentions towards the CE, including preferences for remanufactured products and circularly packaged products [47, 48]. Furthermore, the literature analysed companies’ perspectives on the transition to a circular economy, how they incorporate circular economy initiatives into their production operations, and the obstacles they face in this process [49, 50, 51]. To ensure broader coverage of CE-related news data, 53 keywords were included in the experiment to ensure Guardian news data can be obtained more widely.
Digitalisation has significantly reshaped firm’s production models in recent years [52], extending its influence beyond information technology to various aspects, including the circular practices. For instance, digital technologies play a crucial role in achieving carbon neutrality by enhancing product value and reducing pollution [53]. The application of digital technology not only improves the efficiency of the CE but also allows companies to broaden the scope of its applications and effectively utilise scarce resources [54]. Furthermore, integrating the circular economy concept into product and service design substantially enhances consumption potential by decoupling growth from resource availability [48]. Previous research has identified a positive relationship between digitalisation and the development of the CE [54]. Tools such as big data analytics and the Internet of Things (IoT), when combined with digital technology, enable monitoring and tracking of product life cycle data, thus minimising waste and enhancing recycling rates [55]. Artificial intelligence enhances productivity and promotes CE development through optimised processes and real-time data analysis [54]. Additionally, machine learning anticipates uncertainties in the design and production process, identifying deficiencies in the CE system through real-time monitoring [56]. Therefore, analysing the circular economy from a digital perspective can assist companies in achieving a low-carbon transformation and effectively capturing public views and attitudes towards the circular economy.
3 Methodology
3.1 Framework
Figure 2 describes the proposed framework for the interdisciplinary study, with a particular emphasis on investigating trends and concerns within public opinion regarding the CE. Firstly, we collected textual posts from Twitter, Reddit, and The Guardian platforms to reflect the opinions and attitudes of three distinct strata of the public. Then, it is worth to note that to ensure compliance with data regulations, we utilised official APIs to retrieve the relevant data. Next, we conducted exploratory data analysis with the purpose of gaining a preliminary understanding and getting to know the data landscape before taking necessary data preprocessing, i.e., data cleaning. In Figure 2, the red dashed rectangle contains the processes related to data collection, which will be discussed in Section 3.2. Meanwhile, the green dashed rectangle includes hyperparameter optimisation for topic modelling, which will be discussed in Section 3.3.
Figure 2: The proposed multidisciplinary framework for using hyperparameter optimisation on topic modelling in the context of circular economy.
3.2 Datasets
CE is a broad topic that involves different groups of people, such as the public, professionals, and officials. Analysing the various opinions and interests in CE across different groups requires us to collect data from a wide spectrum of society. To achieve this goal, we collected textual posts from Twitter and Reddit, respectively representing the opinions of the public and professionals. Meanwhile, we also scraped The Guardian news to analyse official opinions on CE.
The process of data scraping for this research can be summarised as the following steps:
  1. 1.
    Search suitable websites containing data from various segments of the public individually. Ensure a comprehensive understanding of diverse perspectives.
  2. 2.
    Check the policy of the selected websites to ensure that data scraping is permitted for research purposes and that the process complies with their regulations.
  3. 3.
    Follow the university’s rules to submit ethic approval application.
  4. 4.
    Investigate the usage of the APIs offered by websites. Data scraping can be challenging, especially for websites that reserve large amounts of data, as searching for user-customised data is time-consuming. Additionally, different websites may design different APIs with varying arguments (e.g., keywords, time span) to support efficient data scraping. Therefore, it is important to be familiar with the API usage to ensure that the downloaded data meets expectations.
In machine learning, data preprocessing is a critical step for models to achieve satisfactory performance. This step is also vital for topic modelling because we have found that the quality of the original scraped data is poor and not suitable for direct use in topic models. For example, website links commonly appear in social media posts and news; however, they are less informative for topic modelling and may adversely affect model performance. In this research, data preprocessing primarily involves several targets:
  1. 1.
    Stopwords removal
  2. 2.
    Punctuation removal
  3. 3.
    Email and URL removal
  4. 4.
    Special symbols removal
  5. 5.
    Lemmatisation
  6. 6.
    Check whether the text length exceeds the model limit
3.3 Hyperparameter Optimisation for Topic Modelling
Hyperparameters in machine learning denote the parameters that need to be set before the model is trained or fitted. In other words, a machine learning algorithm can be instantiated by configuring it with a set of hyperparameter values. These values control the behaviour of algorithms and also affect aspects such as model complexity and convergence. Therefore, different sets of hyperparameter values can produce different results.
Hyperparameter selection inherently is an iterative process of trial and error, and the selection usually relies on a quantitative metric. In our research, to compare different sets of hyperparameter values, we used the coherence (NPMI) value, which ranges from -1 to 1; a value of 1 indicates the best coherence. Manual tuning of hyperparameters is very time-consuming and relies on machine learning intuition and experience. Therefore, many advanced approaches have been proposed to automate the process of hyperparameter optimisation, such as CMA-ES [57] and TPE [58]. However, these approaches usually perform well and exhibit their strengths on relatively large hyperparameter search spaces.
In this study, we selected hyperparameters to be optimised for three models according to the suggestions from original papers: BERTopic [27], CorEx [12], and LDA [42]. We have decided to collaboratively use BERTopic, CorEx, and LDA for topic modelling experiments in our research, based on their complementary analytical strengths, which together provide a robust methodological framework for dissecting complex changes in public sentiment. BERTopic leverages the powerful capabilities of large language model architectures, enabling it to handle vast amounts of unstructured text with a high degree of semantic understanding. This allows BERTopic to extract thematic clusters that reflect intricate details of public sentiment on the CE. CorEx employs an information-theoretic approach, capable of identifying subtle patterns and correlations within the data [12]. This makes the CorEx model particularly valuable in parsing the thematic structures behind public discourse, thus complementing the depth of BERTopic through meticulous detection of latent topics. LDA offers a probabilistic perspective on topic modelling, providing a broad overview of the topic distributions within the data. LDA’s strength lies in its ability to generalise across large texts, thus enabling a macro-level understanding of the popularity and shifts concerning CE topics. These three methods ensure a comprehensive exploration of topic distributions, capture the granularity and breadth of public discourse, and offer a multifaceted understanding of public sentiment towards CE on different online platforms.
Furthermore, the respective search space is defined with the following considerations: (1). Try to make the search space size of the three models equal. (2). We used the model’s default or suggested values as central values to create search spaces by consecutively adding or subtracting a certain number of fixed values. (3). We made a few adjustments for some special cases to refine the search space. For example, in BERTopic, ‘n_cluster’ denotes the number of topics that a model needs to generate. This search space starts from 2 because we consider 2 to be the smallest number of topics. The hyperparameters and search spaces are summarised in Table 1, and each model used the same search space for three datasets.
It is noted that random search and grid search are two popular approaches because they are simple and easy to implement. As shown in Table 1, our defined search spaces are not large compared to those in deep-learning research. Additionally, the three models have relatively lower computational complexities compared to large deep learning models. Therefore, we decided to use grid search to perform an exhaustive search to confirm the best hyperparameters, as using random search in such small search spaces does not make sense.
Moreover, it is worth drawing attention to the fact that topic modelling is a type of unsupervised learning. No single quantitative metric can comprehensively evaluate the results, pointing out the importance of human evaluation. Therefore, in our research, we will begin by comparing the hyperparameter optimisation results in Section 4.1 with human interventions. Subsequently, we will utilise the selected hyperparameter optimisation results to conduct further analysis (with visualisation) in the context of circular economy in Sections 4.2 $\sim$ 4.4.
Hyperparameter Search Space
BERTopic n_gram [1,3]
n_clusters [2, 5, 10, 15, 20, 25]
n_components [5, 10, 15]
n_neighbors [10, 15, 20]
CorEx anchor_strength [1, 1.5, 2, 2.5, 3, 3.5, 4, 4.5, 5, 5.5]
n_hidden [3, 5, 10, 15, 20, 25]
LDA n_topics [2, 5, 10, 15, 20, 25]
alpha [0.04, 0.05, 0.07, 0.1, 0.2, 0.5]
eta [0.04, 0.05, 0.07, 0.1, 0.2, 0.5]
Table 1: The hyperparameters selected to be optimised and their corresponding search spaces.
4 Experiments and results
4.1 Hyperparameter Optimisation
The results of hyperparameter optimisation for three selected models (BERTopic, CorEx, and LDA) on three datasets (The Guardian, Reddit, and Twitter) are shown in Table 2. The coherence values (NPMI) presented in Table 2 indicate that BERTopic outperformed the other two models on two of three datasets. However, much of the research in topic modelling emphasises that evaluating the quality of topic modelling should not solely rely on quantitative metrics but also incorporate human evaluation. Therefore, we listed the topics obtained from three topic models with the best hyperparameters on three datasets in Sections 4.1.1 $\sim$ 4.1.3, respectively. In the following sections, we will compare, analyse, and discuss the topics obtained from different models, dataset by dataset, to decide which model will finally be selected for further analysis.
Model Dataset Best Parameters Best Coherence (c_npmi)
BERTopic The Guardian
n_gram=2.0,
n_clusters=2,
n_components=5,
n_neighbors=20
0.1669
Reddit
n_gram=2.0,
n_clusters=2,
n_components=15,
n_neighbors=20
-0.0609
Twitter
n_gram=2.0,
n_clusters=10,
n_components=10,
n_neighbors=10
0.1354
CoRex The Guardian
anchor_strength=5.5,
n_hidden=20
0.1244
Reddit
anchor_strength=1.5,
n_hidden=3
-0.1428
Twitter
anchor_strength=1.0,
n_hidden=5
0.1538
LDA The Guardian
n_topics=5.0,
alpha=0.5,
eta=0.5
-0.0451
Reddit
n_topics=2.0,
alpha=0.5,
eta=0.04
-0.0671
Twitter
n_topics=2.0,
alpha=0.5,
eta=0.2
0.0194
Table 2: Best hyperparameter values and corresponding coherence scores for BERTopic, CorEx, and LDA models on three datasets.
Topic Topic Term
0 energy, gas, uk, fuel, carbon, emission, increase, power, plan, rise
1 australia, australian, nsw, minister, leader, labor, continue, say, question, need
Table 3: The topic terms over 2 topics obtained from the BERTopic model with the best hyperparameter values on the Guardian dataset.
Topic Topic Term
0 reduce, emission, energy, carbon, gas, climate, renewable, fuel, fossil, greenhouse
1 reuse, design, fashion, clothe, designer, brand, shop, buy, glass, kitchen
2 recycle, plastic, waste, recycling, landfill, product, food, material, packaging, bottle
3 year, time, come, photograph, day, week, month, tell, high, early
4 minister, election, party, leader, prime, vote, parliament, political, liberal, speak
5 water, environmental, specie, environment, forest, wildlife, natural, land, biodiversity, river
6 update, bst, gmt, australian, australia, today, morning, labor, albanese, aap
7 police, military, russian, war, court, crime, russia, criminal, ukraine, attack
8 covid, 19, coronavirus, virus, vaccine, pandemic, infection, lockdown, outbreak, vaccination
9 idea, century, society, kind, sense, culture, history, inspire, modern, bear
10 man, love, old, game, fan, star, player, play, season, ball
11 think, thing, like, know, look, want, ask, feel, way, good
12 government, report, say, increase, policy, state, support, public, plan, new
13 price, economy, inflation, tax, cost, economic, rise, bank, financial, growth
14 company, industry, market, business, investment, technology, consumer, sector, sustainable, sustainability
15 trump, biden, republican, donald, president, joe, american, administration, washington, americans
16 health, child, people, school, hospital, death, care, woman, family, medical
17 tory, labour, conservative, uk, secretary, boris, sunak, johnson, rishi, chancellor
18 country, world, global, international, china, africa, germany, brazil, india, summit
19 city, local, resident, town, road, mile, park, area, car, travel
Table 4: The topic terms over 20 topics obtained from the CorEx model with the best hyperparameter values on the Guardian dataset.
Topic Topic Term
0 say, water, year, people, health, new, area, city, environmental, environment
1 year, like, £, time, photograph, car, good, go, day, new
2 say, government, energy, climate, year, uk, new, emission, £, change
3 say, bst, gmt, people, update, government, minister, australia, new, party
4 work, people, say, year, business, company, need, like, way, use
Table 5: The topic terms over 5 topics obtained from the LDA model with the best hyperparameter values on the Guardian dataset.
4.1.1 The Guardian
In Table 2, BERTopic achieved the best performance with a coherence value of 0.1669. To further investigate the performances of the three models, Tables 3 to 5 show the topics generated by three different models on the Guardian dataset. From our perspective, LDA topics are not desirable because the word ‘say’ appears and dominates in its three topics, but it is less meaningful for CE. Meanwhile, BERTopic and CorEx are relatively good; we can clearly see words such as ‘energy’, ‘gas’, and ‘power’, which are closely related to CE. Although we found ‘Australian’-related words in the topics, we manually checked our dataset and found that there are some news articles from the Australian news sector scraped by the API. Due to the large size of the dataset, we are not able to delete them manually, so we kept them and take the advantages of topic modelling to ignore them in the following analysis.
number value params_n_clusters params_n_components params_n_gram params_n_neighbors
1 0.16694393 2 5 2 20
2 0.16458788 2 5 3 10
3 0.16329703 5 15 3 15
4 0.14380017 20 15 1 10
5 0.13858814 2 5 3 20
Table 6: The top 5 hyperparameter sets for BERTopic on the Guardian dataset
However, there are only two topics generated by BERTopic, which is less informative for our subsequent analysis. Therefore, we revisited the BERTopic hyperparameter optimisation results, expecting to find a set of hyperparameter values that can achieve a relatively high coherence value and produce informative topics at the same time. We list the top 5 hyperparameter sets in Table 6 obtained by hyperparameter optimisation (HPO) on BERTopic, referring to coherence values. We found that the number 4 hyperparameter set in Table 6 has the same number of topics as CorEx, and the coherence value 0.14380017 is higher than that of CorEx (0.1244). Therefore, we selected the hyperparameters of BERTopic from the number 4 hyperparameter set for our subsequent analysis. The new topics are shown in Table 7.
Topic Topic Term
0 australia, australian, nsw, election, labor, bst, senator, coalition, need, leader
1 renewable, carbon, energy, emission, international, environment, world, resource, goal, climate
2 energy, renewable, uk, gas, fuel, carbon, solar, britain, emission, electricity
3 mp, corbyn, eu, britain, tory, uk, brexit, labour, conservative, british
4 recycling, recycle, recyclable, plastic, waste, reuse, circular, packaging, environment, environmental
5 sustainable, environment, farming, green, land, grow, plan, uk, farmer, produce
6 sustainability, sustainable, environment, approach, resource, ethical, consumer, environmental, growth, future
7 australia, energy, renewable, nsw, australian, 2050, emission, gas, fuel, carbon
8 candidate, biden, trump, election, voter, republicans, official, democrat, campaign, senator
9 vaccine, lockdown, covid, vaccination, outbreak, vaccinate, quarantine, coronavirus, gmt, health
10 ftse, eurozone, gmt, trading, inflation, market, uk, price, investor, european
11 feel, moment, want, people, audience, picture, say, tell, die, like
12 airbnb, rent, sharing, share, create, offer, site, host, idea, uber
13 player, sport, goal, liverpool, league, kick, chelsea, want, arsenal, club
14 ev, tesla, car, electric, charge, nissan, vehicle, emission, petrol, driver
15 epa, pfas, chemical, toxic, federal, environmental, lead, pesticide, emission, official
16 ukraine, russia, ukrainian, putin, kremlin, eu, world, russian, nato, kyiv
17 samsung, smartphone, tablet, android, device, price, app, screen, ipad, fold
18 gambling, gamble, betting, gambler, bet, casino, bookie, punter, bookmaker, gaming
19 pga, golfer, golf, fairway, tournament, mickelson, leaderboard, spieth, championship, shoot
Table 7: The topic terms over 20 topics obtained from the BERTopic model with the alternative hyperparameter values on the Guardian dataset.
4.1.2 Reddit
To explore discussions related to the circular economy, we utilised the Reddit API to collect relevant posts for our topic modelling. The Reddit Official API facilitates the extraction of posts by subreddit (for a particular topic) or keyword searches. We specifically targeted subreddits related to the circular economy and filtered for posts containing the keyword ‘circular economy’ across broader discussions. This approach enabled us to compile thousands of posts, including their titles, contents, and time. We store all data in JSON files to ensure that the data format is consistent across platforms.
The BERTopic model (Table 8) identifies two topics. The first topic integrates core aspects of the circular economy such as ‘recycling’, ‘sustainable’ and ‘waste’ pointing towards operational and environmental concerns. The second topic seems to focus on the business and ideational perspectives, including terms like ‘business model’ and ‘circularity’, which suggest strategic to circular economy practices. BERTopic provides a better understanding of the contextual relationships between words compared to traditional LDA models because of the presence of the large language model BERT. By utilising BERT’s deep semantic understanding, it can link concepts such as environment and sustainability, generating related words such as ‘circular economy’ and ‘sustainable’. The word embedding of the BERT model allows the model to capture the link between specific activities such as ‘recycling’ and ‘plastic waste’ and the broader topic of circular economy.
The CorEx model (Table 9) results show a diverse set of subtopics over three topics. The first topic is focused on fundamental actions and strategies like ‘reduce’ and ‘consumption’. The second topic is more oriented towards community and engagement with terms like ‘help’, ‘thank’ and ‘people’. The third topic captures industrial and infrastructural elements, with terms such as ‘material’, ‘battery’ and ‘recycling’ reflecting specific operational sectors. The best hyperparameter results are more comprehensive than BERTopic because CorEx focuses more on discovering the greatest common information between variables, not just combinations of words that occur frequently in LDA. The anchor vocabulary of 3R Rules that we set up in our experiments to guide the model to focus more on specific thematic directions also helped CorEx to identify non-intuitive topics in Reddit. For example, Topic 3 shows that words such as ‘material’, ‘battery’ and ‘recycling’ reflect industry-specific discussions. We believe this provides valuable insight for industry policymakers.
Finally, the LDA model (Table 10) produces two topics that, while similar, emphasise different facets of circular economy discussions. Both topics feature ‘circular’, ‘economy’ and ‘waste’ as central terms but differ slightly in focus. The first topic includes ‘new’ and ‘need’, which may indicate discussions around solutions within the circular economy. The second topic includes ‘business’ and ‘good’ likely reflecting on the efficacy and benefits of circular economy practices in business contexts. From the words frequently appearing in the two topics, such as ‘economy’, ‘circular’, ‘waste’ and ‘sustainable’, it can be seen that the LDA model directly captures the core words related to circular economy. While this approach may not be as sensitive as deep learning-based BERTopic, it has a unique advantage in providing intuitive topic distribution.
Topic Topic Term
0 circular_economy, recycling, recycle, plastic waste, circular, sustainable, reuse, environment, sustainability, waste
1 circular_economy, circular, business_model, idea, sustainable, recycling, sustainability, resource, circularity, recycle
Table 8: The topic terms over 2 topics obtained from the BERTopic model with the best hyperparameter values on the Reddit dataset.
Topic Topic Term
0 reduce, consumption, waste, world, enable, city, use, base, produce, bag
1 help, thank, work, project, reuse, hi, look, people, like, currently
2 recycle, material, battery, company, loop, london, recycling, solar, public, clean
Table 9: The topic terms over 3 topics obtained from the CorEx model with the best hyperparameter values on the Reddit dataset.
Topic Topic Term
0 economy, circular, new, waste, product, plastic, help, recycle, sustainable, need
1 circular, economy, waste, plastic, business, sustainable, product, new, good, help
Table 10: The topic terms over 2 topics obtained from the LDA model with the best hyperparameter values on the Reddit dataset.
4.1.3 Twitter
While exploring user sentiments on the CE from Twitter, we utilised Twitter’s official API to collect posts containing the hashtag ‘Circular Economy’. However, starting in 2023, Twitter’s official API began to restrict the scope of user searches, allowing us only to access data from the past year. We conducted data cleansing on all collected data, retaining core information such as the content and timing of tweets. We also removed duplicates, replied and reposted tweets, ultimately preserving 3,922 English tweets for our model training.
Table 11 shows topics from the Twitter dataset by the BERTopic model. Topic 0 clusters around the principles of the circular economy, emphasising the integration of recycling and waste management into sustainable practices. The convergence of ‘sustainable’ and ‘circular’ within this topic highlights a prevailing narrative focused on the implementation of resource-efficient economies. Topic 1 projects a future-oriented discussion, where the aspiration for a ‘sustainable future’ is intricately linked to the notion of ‘circularity’, reflecting the community’s advocacy for systemic change towards enduring environmental stewardship.
Furthermore, Topic 2 focuses on the recyclability of packaging materials, illustrating a deepened public concern over the lifecycle of consumables and the pursuit of sustainable resource utilisation. In an interesting thematic diversion, Topic 3 intertwines charity-related activities with the circular economy, suggesting that philanthropic efforts are increasingly oriented towards supporting sustainable practices. Topics 5 through 9 pivot to a technological perspective, shedding light on discussions around the resale and recycling of network equipment. The presence of terms specific to networking hardware such as ‘router’ and ‘Cisco’ points to a vibrant discourse on the CE within the tech sector.
Topic Topic Term
0
circular economy, sustainability circulareconomy, circulareconomy sustainability, recycling,circulareconomy, recycle,wastemanagement, circular, sustainable,
sustainability,
1
circulareconomy, sustainability circulareconomy, circulareconomy sustainability, circulareconomy circulareconomy, circular economy, sustainable future, circularity,
circular, create sustainable, sustainable,
2
sustainablepackage recyclability,recyclability package,circulareconomy recycledmaterial,sustainablepackage,recycledmaterial resourceefficiency,
package sustainability,recycle sustainability,sustainability circulareconomy,circulareconomy sustainability,circulareconomy recycle,
3
monday charitynee,supportlocal mondaythought,charity need,need charity,charity circulareconomy,work charity,support need,need monday,support amazing,
help support,
4
ban plastic,plasticwaste export,circulareconomy endwastecolonialism,plastic waste,endwastecolonialism tell,wasteshipment breakfreefromplastic,
mess wastetrade,waste export,petition wastetrade,wastetrade,
5 router processor,processor cisco,cisco 12000,performace router,router cisco,forsale cisco,cisco circulareconomy,2951 router,cisco 2921,asr1001 router,
6 forsale cisco,c3650 48td,ws c3650,c3650,switch 3650,48 port,cisco,cisco ws,cisco circulareconomy,48td catalyst,
7 10xge port,adapter xfp,xfp port,spa 1xtenge,cisco spa,port adapter,gigabit ethernet,port multifunction,forsale cisco,1xtenge xfp,
8 extreme network,24pt poe,switch extreme,summit x450e,summit x460,summit x450a,poe switch,cable extreme,x450e 24p,x460 24,
9 roast cup,roaster chef,roaster,coffee roaster,mb roast,roast,roaster barista,contemporary roaster,cheflife chefsofinstagram,cup chef,
Table 11: The topic terms over 10 topics obtained from the BERTopic model with the best hyperparameter values on the Twitter dataset.
As show in Table 12, the CorEx model’s analysis of the Twitter dataset reveals a nuanced landscape of sustainability discourse. Topic 0 focuses on reducing environmental impacts, emphasising emissions, sustainable resource usage, and the need for carbon footprint minimization. Topic 1 reflects community engagement in sustainability through reuse and repair, underscored by charitable activities. Topic 2’s dialogue centres on recycling specifics, indicating targeted discussions on material recovery and recycling industries. Topic 3 diverges to a commercial domain, highlighting events and technology sales, possibly in an ecological context. Finally, Topic 4 focused on global waste management policies.
Topic Topic Term
0 reduce, emission, carbon, resource, waste, sustainable, planet, recyclability, sustainablepackage, co2
1 reuse, charity, item, repair, charitynee, monday, supportlocal, amazing, mondaythought, volunteer
2 recycle, recycling, tyre, tyrerecycle, battery, tirerecycle, tyrerecovery, recyclingbusiness, noexcusesgocircular, tire
3 forsale, register, event, cisco, april, conference, 2023, port, summit, join
4 breakfreefromplastic, wastetrade, baselconvention, wasteshipment, ban, endwastecolonialism, export, mess, tidy, eu
Table 12: The topic terms over 5 topics obtained from the CorEx model with the best hyperparameter values on the Twitter dataset.
Finally, according to the topics generated by the LDA model on the Twitter data set (Table 13): Topic 0 encapsulates the environmental concerns associated with the circular economy, with a strong focus on ‘waste’, ‘recycle’, and ‘plastic’. The inclusion of ‘eu’, ‘breakfreefromplastic’, and ‘wastetrade’ within this topic suggests a geographically and politically contextualised discussion. In Topic 1, terms like ‘need’, ‘waste’, and ‘package’ alongside ‘sustainablepackage’ indicate an exploration of how businesses can address sustainability through innovative packaging solutions.
In conclusion, BERTopic provided a deep extraction of CE topics through a large language model, CorEx provided a unique separation of themes (highest coherence score), and LDA captured the underlying concepts of CE on Twitter, all of which provide valuable insight into our understanding of the public sentiment on the CE. In the next three subsections, we will visualise and discuss the topics generated by the best models on each dataset in more depth.
Topic Topic Term
0 circulareconomy, waste, recycle, plastic, sustainability, new, eu, breakfreefromplastic, wastetrade, wasteshipment
1 circulareconomy, sustainability, circular, economy, sustainable, business, need, waste, package, sustainablepackage
Table 13: The topic terms over 2 topics obtained from the LDA model with the best hyperparameter values on the Twitter dataset.
4.2 Official Sources: The Guardian
As illustrated in Figure 3, we have selected seven topics from the BERTopic model applied to The Guardian dataset for analysis. The bar chart displays the scoring (c-TF-IDF score) of individual words (terms) within each topic, ranging from 0 to 1. Higher scores indicate a greater significance of the word in its respective topic, reflecting its substantial influence on the topic’s importance.
It is evident that under optimal parameter tuning, the BERTopic model yields scores of 0.6 or higher for every word within the seven selected topics. Topics 1 and 5 demonstrate the most uniform distribution of word scores among the ten words analysed, indicating a lack of distinct representative words that characterise these topics compared to others. The majority of words in Topic 1 pertain to global aspects of energy and the environment, with ‘renewable’ achieving the highest score and ‘climate’ the lowest. In contrast, Topic 5 does not feature words related to ‘world’; instead, it includes terms such as ‘UK’ and ‘farmer’, focusing more on national-level sustainable agricultural development.
The remaining five topics each feature one or two words with significantly higher scores. For instance, in Topic 14, the score for ‘ev’ is notably higher than that of the other words. This is indicative of the strength of BERTopic as a large language model that utilises contextual analysis to effectively summarise and comprehend abbreviations like ‘ev’. Other related terms, such as brands including ‘Tesla’ and ‘Nissan’, aid the model in more swiftly grasping the significance of ‘ev’. The descending order of scores also illustrates the prioritisation and segmentation of these words within the topic.
We acknowledge that BERTopic performs admirably in interpreting text within The Guardian dataset, aggregating relevant vocabulary under a unified topic. However, an analysis of the highest-scoring words from Topics 4 and 6 reveals that words of different grammatical forms often convey essentially the same meaning. Both ‘recycling’ and ‘recycle’ in the text reflect the concept of resource recovery, while ‘sustainability’ and ‘sustainable’ repeatedly express the notion of environmental sustainability. We believe that by leveraging the strengths of large language models, including part-of-speech normalisation and selective word comparison, BERTopic can offer researchers a more comprehensive vocabulary pertaining to each topic.
Figure 3: Bar chart representation of topic modelling results using the BERTopic from The Guardian dataset.
The frequencies and corresponding times of the seven themes are presented in Figure 4. Topic 1 shows a growing trajectory, rising significantly starting around 2017. This upgrade is timed to coincide with key international developments such as the 2016 Paris Agreement, highlighting the growing focus in the UK news on the synergies between renewable energy adoption and carbon emissions. Topic 15 experienced a surge in frequency around 2019. The topic continues to gain visibility, and this rise may be related to increased public concern about chemical pollution incidents and environmental health. The increase in frequencies of all topics post-2015 suggests an overall growing public and journalistic interest in environmental issues. This period seems to be pivotal in shaping the public sentiment on sustainability and environmental topics.
Figure 4: This line chart shows the frequency of occurrence of seven topics generated by the BERTopic model trained on The Guardian dataset (1999 to 2023).
Figure 5 presents a heatmap of the similarity matrix for topics generated by the BERTopic model on The Guardian dataset. These seven topics correspond to the ones shown in Figure 3. Deeper shades represent higher cosine similarity, whereas lighter shades indicate lower similarity. Since we get more topics in the Guardian dataset than in the Reddit and Twitter dataset, we believe the heatmap in Figure 5 facilitates a comparative analysis of the interrelationships among the topics.
The relationship between Topic 2 and Topic 15 is indicated by the darker shades on the heatmap, suggesting a strong connection. Further investigation into the vocabulary reveals that discussions on renewable energy in The Guardian are considered within a broader context, where the impact of chemical regulations and substances such as PFAS is indispensable. There could be a policy-oriented discourse that integrates renewable energy initiatives with the regulation of hazardous chemicals. Compared to its similarity with Topic 15, Topic 2 may have a closer relationship with Topic 4. Topic 2 and Topic 4 both involve environmental sustainability. ‘Recycling’ is a key element in Topic 4 and an integral part of sustainable practices. The discussions in Topic 4 likely address practical issues of environmental sustainability, such as ‘waste’ and ‘packaging’ which complement the broader topic of sustainability in Topic 2.
The relatively superficial relationship between Topics 5 and Topic 6 is particularly intriguing, given that both seemingly focus on sustainability and the environment. However, the low similarity score indicates that each topic explores different facets of this broad thematic focus. Topic 5 encompasses terms such as ‘sustainable’, ‘environment’, ‘UK’, ‘farmer’, and ‘produce’. These terms collectively suggest discussions centred around sustainable agriculture and production. ‘Farmer’, ‘farming’, and ‘produce’ highlight the practical implementation of sustainable practices within agricultural activities under the CE. The presence of ‘green’ and ‘land’ signifies concerns over ecological considerations in land use. The terms ‘plan’ and ‘UK’ may denote specific strategies within the British agricultural sector. Conversely, Topic 6 is characterised by words like ‘sustainability’, ‘sustainable’, ‘approach’, ‘resource’, ‘ethical’, ‘growth’, and ‘future’. While both topics emphasise sustainability and the environment, the additional presence of terms such as ‘approach’, ‘resource’, and ‘ethical’ indicates a broader conceptual discourse. The term ‘ethical’ implies discussions on the moral imperatives of sustainability, potentially involving corporate responsibility, while ‘consumer’ shifts the focus to individuals and the market. ‘growth’ and ‘future’ suggest a forward-looking perspective, contemplating the long-term trajectory of sustainable development. Although there is thematic overlap in the general concern for sustainability and environmental issues, BERTopic effectively reveals the specific contexts in which these terms are used, highlighting divergences in discourse. Topic 5 focuses more on the practicalities of sustainable agriculture and environmental management. In contrast, Topic 6 appears to discuss sustainability from a broader perspective, encompassing ethical considerations, consumer habits, and approaches aimed at promoting long-term sustainable growth.
Further inspection of the heatmap we obtained, we found topic 14 and topic 15, which have less overlap with other topics. The lighter shading indicates this visually, suggesting that the discourse within these topics is more isolated and specialised. Topic 14’s focus on electric vehicles also suggests that the Guardian’s discussion, largely represented by Tesla, opens up a dialogue with the wider environment. Likewise, Topic 15 deals with chemical regulations, with a possible emphasis on PFAS and EPA oversight, suggesting that the focus of the discussion is on the specificity of regulation and science and does not broadly intersect with the rest of the topic landscape. Despite the differing focuses of Topics 14 and Topic 15, we observed that their substantial similarity may stem from a shared concern for environmental impact and regulation. Both topics likely engage in a broader narrative about reducing harmful emissions, whether through the promotion of electric vehicles or the regulation of toxic chemicals. Furthermore, both topics appear to involve human activities (driving in Topic 14 and chemical usage in Topic 15) and governmental measures to mitigate associated risks. This convergence may reflect an increasing societal and regulatory discourse focused on sustainability and mitigating environmental hazards.
Figure 5: Heatmap of the similarity matrix between topics from The Guardian dataset using the BERTopic.
Figure 6 provides a two-dimensional representation of the topics generated under BERTopic for The Guardian dataset, with each data point representing an individual document. The position of each document in two-dimensional space (depicted by D1 and D2) reflects the topic of the document as determined by BERTopic’s semantic processing. A colour-coding scheme is employed to demarcate document clusters that correspond to the discrete topics unearthed by BERTopic.
According to this scatter plot, we can notice that the cluster of Topic 14 (pink colour) is somewhat separated from the rest of the topic clusters. This means that the news around electric vehicles in The Guardian is more of a dedicated discussion to this area, focusing on key industry players such as Tesla. While this topic may not be at the centre of most discussions in The Guardian news, it can be a good representation of the impact of innovation trends, CE being carried out in transport. Additionally, although the Topic 15 (grey colour) is similar to other clusters, it has a smaller size which may reflect the specificity of discussions around toxic chemicals. The size of clusters in this topic shows there is a specialised and smaller number of news about chemical term and EPA in The Guardian dataset.
Figure 6: Scatter plot visualisation of documents and their associated topics from The Guardian dataset, clustered by similarity via the BERTopic model.
In the field of topic modelling, word clouds serve as a visual heuristic method, displaying terms that occupy central nodes in the discourse network [59]. Figure 7 shows the topic vocabulary graph cloud generated by BERTopic on The Guardian dataset. The visual prominence of terms such as ‘sustainability’, ‘sustainable’, ‘recycling’, and ‘energy’ within the cloud is indicative of their pivotal role in the corpus, highlighting the dataset’s focus on environmental conservation, renewable energy, and the principles of the CE. These terms underscoring the centrality of sustainable energy in CE are captured within The Guardian news. Surrounding these core terms are secondary, yet substantial, concepts such as ‘ev’, ‘electric’, and ‘vehicle’, which, given their relative visual weight (size), suggest substantive topics that concern technological advancements in transportation and the transition. Moreover, the adjacency of concepts such as ‘ethical’, ‘consumer’, ‘growth’, and ‘future’ alongside ’sustainability’ within the visualisation elucidates an interplay of notions.
Figure 7: Word cloud visualising the prominence of terms in topics derived from The Guardian dataset using the BERTopic model.
4.3 Professional Sources: Reddit
Figure 8 illustrates all topics generated by the BERTopic model on the Reddit dataset using the best hyperparameters. Through hyperparameter optimisation, we determined that configuring the generation of two topics yields optimal results on the Reddit dataset. Hence, the diagram displays only these two topics alongside the scoring of various words. The word scores of Figure 8 (0 to 1), adhering to the same criteria as illustrated in Figure 3, serve as an indicator of the contextual relevance of words within a topic. This provides an empirical foundation for the analysis of topics generated by the BERTopic model on the Reddit dataset.
Unlike the results from The Guardian dataset, the topics generated on Reddit by the BERTopic model include not only individual words but also phrases composed of two words. In Topic 0, terms such as ‘circular economy’, ‘recycling’ and ‘plastic waste’ dominate, indicating that the overall narrative focuses on environmental management, particularly within the scope of waste management. The recurrent topic of ‘circular economy’ supports discussions that emphasise material reuse as a strategy for reducing waste production.
Conversely, Topic 1 encapsulates vocabularies that link environmental issues with business. Words such as ‘business model’, ‘sustainable’ and ‘resource circularity’ substantiate these linkage. Topic 1 points towards the concept of sustainable development, which is integral to business innovation, combining economic incentives with ecological responsibilities. The repeated occurrence of words like ‘circular economy’ and ‘recycling’ across both topics underscores their central role in the discourse. However, it can be observed in Topic 1 which is generated by the BERTopic model on the Reddit dataset, the emphasis on ‘business model’ and ‘idea’ suggests a more strategic perspective than the operational viewpoint found in Topic 0.
Figure 8: Bar chart representation of topic modelling results using the BERTopic from Reddit dataset.
In Figure 9, Topic 0 involves terms such as ‘circular economy’, ‘recycling’, ‘recycle’, ‘plastic waste’, and related environmental terms. After 2015, it shows a declining trend, stabilising in subsequent years with fluctuations but a continuous presence. This trend may indicate that discussion peaks on Reddit could have been triggered by specific events or policies. Each fluctuation could reflect the evolving interest in the circular economy (CE) on Reddit. As concepts of waste management and sustainable practices increasingly penetrate collective consciousness, they might also represent an evolving discourse. Topic 1 (yellow line) focuses on ‘circular economy’, ‘circular’, ‘business model’, and ‘idea’. Topic 1 displays greater volatility and pronounced peaks, particularly in 2022, which may relate to significant international decisions on climate and economy or the easing of the COVID-19 pandemic. Topic 1 shows more distinct peak trends, while Topic 0 exhibits more consistency after an initial decline. This may suggest that discussions on Reddit about circular economy business models are more sensitive to specific events and also reflect ongoing discussions within the community about the development of CE.
Figure 9: This line chart shows the frequency of occurrence of the two Topics generated by the BERTopic model trained from the Reddit dataset in recent years.
Figure 10 reveals the document distribution and description of two topics as a scatter plot. Topic 0 is shown in orange and summarises the discussion with ‘circular economy’ and ‘recycling’. Compared with the distribution of Topic 1, it can be found that the data points are relatively denser. It shows that the conversation space on Reddit is more concentrated, and users’ conversations on Reddit are more focused on the views between CE and recycling. The green Topic 1 discussion revolves around ‘circular economy’ and ‘business model’. It shows that the conversation space on Reddit is wider, and users can have a wider discussion about integrating sustainability into business models. Although these two clusters have different topic focuses, there are still some data points that are close to each other and another topic cluster, reflecting a certain dependence between recycling and business models.
Figure 10: Scatter plot visualisation of documents and their associated topics from Reddit dataset, clustered by similarity via the BERTopic model.
Figure 11 is a word cloud produced by BERTopic on the Reddit dataset, which presents the main topic words of CE in different sizes and positions. Unlike the word cloud (Figure 7) from The Guardian dataset, ‘circular economy’ appears prominently in the word cloud, its large size indicating a central role in the discourse. Adjacent to ‘circular economy’ are terms such as ‘reuse’ and ‘resource’. Although these terms are not as visually prominent as ‘circular economy’, they remain crucial in the narrative, indicating strong discussions about re-usable resources. The central appearance of ‘plastic waste’ marks focused conversations around plastic consumption and waste management. The visibility of this term in the cloud summarises community engagement with specific environmental impacts, resonating with global concerns over plastic pollution. Intersecting with these environmental terms, ‘business model’ anchors the discourse in the economic domain. The combination of ‘business model’ and ‘circular economy’ reflects discussions on business structures aligned with circularity and resource conservation.
The word cloud (Figure 11) not only captures the direction of Reddit user’s discussions about CE but also provides insight into the collective focal points of the community, thereby reflecting the broader public engagement and sentiment towards sustainability.
Figure 11: Word cloud visualisation of the prominence of terms in topics derived from Reddit dataset using the BERTopic model.
4.4 Public Sources: Twitter
This section analyses public participation and perspectives on the circular economy as reflected in the Twitter dataset. Using the CorEx topic model, Figure 12 illustrates public attention and discussion on various aspects of the circular economy. The bar chart distinctly identifies the primary topics, revealing high concern for topics such as sustainability, recycling, and environment. This reflects significant public discourse surrounding sustainability and the circular economy. Specifically, Topic 0 focuses on sustainable practices and awareness, highlighting public attention to sustainable development and emphasising the need for ongoing ecological balance and responsible environmental management. Notably, terms like ‘emission’, ‘reduce’ frequently appear in discussions about the circular economy. The prominence of this term indicates public concern and interest in a low-carbon transition, likely spurred by growing awareness of environmental issues and the impacts of global climate change. Topic 2 centres on recyclability and waste management, showcasing public focus on the processes and benefits of recycling, and emphasising its crucial role in resource conservation and waste reduction. ‘recyclability’ is a pivotal public concern, reflecting the importance of waste management in public discourse. The term’s high frequency can be attributed to the global movement and local regulations that promote recycling as a key component of resource conservation and the circular economy.
Figure 12: Bar chart representation of topic modelling results using the CorEx from Twitter dataset.
The line graph (Figure 13) displays the fluctuation frequencies of five distinct topics generated by the CorEx model based on 2023 Twitter dataset. These topics exhibit significant frequency peaks over time, which may correspond to specific events that triggered Twitter discussions or heightened public sentiment. Topic 2 shows less volatility compared to others but experiences a sharp increase in discussion volume. This likely relates to an event connected with ‘recycle’, temporarily bringing this topic to the forefront on Twitter. Notably, Topic 4 demonstrates a very pronounced peak, exceeding the activity of other topics during the same period. This outlier could represent a major event or breakthrough in sustainability towards the end of March 2023, or an activity that briefly dominated discussions. Figure 13 reflects Twitter’s reactive nature as a platform for public discourse. A temporal analysis of topic frequencies can provide deeper insights into how certain themes gain traction within the social media community, highlighting the interplay between global events and public sentiment.
Figure 13: This line chart shows the frequency of occurrence of Topics generated by the CorEx model trained from the Twitter dataset in 2023.
Further insights are provided by Figure 14, which displays a word cloud visualisation that highlights the frequency and significance of specific terms in public attention. The larger, more prominent words like ‘reduce’, ‘carbon’, ‘waste’, and ‘emission’ reflect a strong focus on environmental impact reduction and the urgency of addressing climate change through tangible actions such as waste management and carbon footprint minimisation. The term ‘sustainable’, adjacent to ‘resource’ and ‘planet’, signifies the overarching goal of these discussions—maintaining the earth’s ecological balance. Additionally, ‘recycle’ and ‘recyclability’ underscore the practical measures being discussed, likely linked to the optimisation of material usage and the lifestyle of products. Through this word cloud (14), we can discern multiple aspects of the CE on Twitter. The topics generated by the CorEx model range from individual to broader societal behaviours, all of which assist researchers in playing a role in shaping a sustainable future within the circular economy.
Figure 14: Word cloud visualisation of the prominence of terms in topics derived from Twitter dataset using the CorEx model.
5 Conclusion, Policy Implications and Future Research Avenue
5.1 Conclusion
In the context of sustainable transformation, this study aims to uncover public sentiment of the CE by analysing data from various sources, including social media platforms Twitter and Reddit, as well as the Guardian. Utilising topic modelling, this paper explores the multi-perspective and changing perceptions of the CE among the public, professionals, and officials. Three topic models(LDA, CorEx, and BERTopic) were employed. The findings of this study reveal the public’s growing concern for environmental issues, highlighting widespread focus on topics such as sustainability, recycling, and CE. Furthermore, the dynamic evolution of these topics over time indicates that public interest in the CE is not only increasing but also diversifying. This reflects a deepening engagement with the CE, influenced by global and regional environmental events. Through systematic experiments and optimal hyperparameter tuning of the models, we have successfully captured shifts in public sentiment and opinions, thereby providing a comprehensive view of public concerns regarding the CE. Overall, this study contributes to both academic and practical understandings of public perceptions in the context of the CE. By mapping the cognitive pathways and concerns of different segments of society, this paper aids policymakers and businesses in adapting strategies and communications to promote a more sustainable future.
5.2 Policy Implications
Our study also offers important policy implications. Firstly, this study highlights the growing public sentiment and concern regarding sustainability, recycling, and the CE. To leverage this trend, policymakers should implement targeted education programs to deepen public knowledge and understanding of the CE [60]. These programs can emphasise the benefits of circular practices, illustrating how individuals can contribute to environmental protection through actions such as recycling and choosing sustainable products [61]. Additionally, community engagement programs can be designed to involve local residents in CE initiatives, thereby fostering public awareness and responsibility for environmental stewardship. It is essential that these policies are tailored for different regions and demographic groups to maximise their reach and impact.
Furthermore, the diverse public sentiment on various aspects of the CE, such as emissions reduction and recycling, demonstrates strong support for businesses that adopt CE practices [45]. Policymakers should consider creating incentives for businesses that integrate CE principles into their operations. For example, companies that reduce waste, use recycled materials, or develop innovative products to facilitate a transition to the CE could receive tax exemptions, subsidies, or other incentives [62]. Additionally, policies can be adjusted to encourage companies to take greater responsibility in advancing the CE, ensuring that businesses operate within a clear regulatory framework that promotes sustainable development.
This study also revealed significant and growing public sentiment regarding recycling and waste management. In response, policymakers should therefore enhance existing regulations to establish a more stringent waste management policies and improved recycling processes [63]. Measures could include setting higher recycling targets for government authorities, imposing stricter controls on industrial waste, and providing incentives for companies that significantly reduce waste production. Expanding waste management infrastructure can also help meet the growing public demand for efficient recycling services [64]. Moreover, policies should encourage the development of new technologies that enhance recycling efficiency and waste reprocessing capabilities, ultimately making the CE more effective and accessible.
5.3 Future Research Avenue
We will refine the granularity of our topic models by utilising more advanced hyperparameter tuning algorithms to optimise models with higher coherence NPMI scores. Furthermore, we will conduct deeper explorations based on the BERTopic model, which has shown optimal performance in both datasets used in this study. This includes investigating multimodal approaches that incorporate textual data as well as visual and auditory content, offering a more comprehensive analysis of public sentiment. This might involve integrating Transformer-based architectures [65] with computer vision techniques to analyse images and videos related to the CE on social media, thereby providing a richer and more thorough understanding of public discourse. This integrated approach not only enhances the precision of sentiment and thematic detection but also captures a broader array of data points for sentiment analysis, including emotional expressions and visual symbols. Longitudinal and cross-platform studies are crucial for observing changes in the emotions of different user groups over time, thereby gaining deeper insights into the dynamic shifts in public engagement with sustainability issues.
Data Statement
All data used in our experiments are from The Guardian, Reddit and Twitter, which are publicly available and can be accessed by APIs in The Guardian Open Platform, Reddit API, and Twitter API.
Acknowledgements
This research is funded by Digital Circular Electrochemical Economy (DCEE) EP/V042432/1, and the UK Research and Innovation (UKRI) Interdisciplinary Centre for Circular Chemical Economy EP/V011863/1.